### v1

# A Goemans-Williamson type algorithm for identifying subcohorts in clinical trials

Abstract
We design an efficient algorithm that outputs a linear classifier for identifying homogeneous subsets (equivalently subcohorts) from large inhomogeneous datasets. Our theoretical contribution is a rounding technique, similar to that of Goemans and Williamson (1994), that approximates the optimal solution of the underlying optimization problem within a factor of $0.82$. As an application, we use our algorithm to design a simple test that can identify homogeneous subcohorts of patients, that are mainly comprised of metastatic cases, from the RNA microarray dataset for breast cancer by Curtis et al. (2012). Furthermore, we also use the test output by the algorithm to systematically identify subcohorts of patients in which statistically significant changes in methylation levels of tumor suppressor genes co-occur with statistically significant changes in nuclear receptor expression. Identifying such homogeneous subcohorts of patients can be useful for the discovery of disease pathways and therapeutics, specific to the subcohort.
1 Introduction
A linear separator in $\mathbb{R}^{d}$ is a basic tool used in statistics and machine learning to partition a set of points into two subsets. Such a separator implicitly encodes information about the pairs of points, which it has to separate, in its coefficients. Given a set of Red and Blue points to be separated, there may not exist a linear classifier that separates most of the Red and Blue points. Moreover, even if one exists, then it can be expected to have most coefficients to be far from zero, i.e., it’s not sparse, unless there are structural assumptions on the dataset. On the other hand, if we want a linear separator that separates a small (but non-trivial) percentage of Red points from Blue points, then we may reasonably expect that a sparse linear separator exists, i.e., one which sets most of its coefficients to zero, and it still has low error. This is because the number of Red-Blue pairs separated is smaller, so the amount of information to be encoded in the linear separator is lower. Such separators also have the advantage that they are explainable – the role and importance of each feature is clear from its coefficient. Therefore, if one assumes the separator is sparse, the focus shifts to identifying a subset of Red points to carve out from the rest, so that: (1) the number of Blue points in the identified subset is relatively small (specificity is high); (2) the relative number of Red points in the subset, compared to the total number of Red points, is large (sensitivity is high); and (3) the number of non-zero coefficients in the separator is small (the separator is sparse). Computing such linear separators are the subject of this paper, and we devise a semidefinite programming (SDP) based algorithm (see Algorithm 1) for that purpose. Furthermore, we prove theoretical guarantees on the size of the identified subset (see Theorem 2.2).
One application for such an algorithm is in identifying homogeneous subsets (equivalently subcohorts) of patients in clinical trials. Diseases like breast cancer, sepsis and even bacterial infections are fairly inhomogeneous – the same disease may have very different underlying causes (different mutations, different strains of bacteria, and so on), and therefore different treatments. No single treatment works for a majority of the patients. Therefore, given a drug that works on a small percentage of patients overall, one recourse is to identify a not too small subset for which the drug treatment works with higher likelihood. Moreover, one wants the identifying test to rely on a handful of features (few genes, few mutations etc) – to ensure explainability of the test, and also control for sequencing costs in the field. This is precisely the abstract problem delineated in the previous paragraph – we are willing to trade sensitivity for specificity. Thus, in this paper, we use the algorithm mentioned above (Algorithm 1), on the RNA microarray dataset for breast cancer from CSC+ (12); PCR+ (16); RSS+ (19) [^1] We use this particular dataset because it is large and balanced, i.e., both metastatic and non-metastatic cases are adequately represented. Moreover, it has epigenetic data as well as RNA expression data, and linking the two is an interesting open question in biology CH (19). to identify subsets of metastatic breast cancer patients that have statistically significant over-expression of one or more nuclear receptors, compared to the full dataset (see Figure 3). It’s worth noting here that nuclear receptors are often chosen as drug targets OALH (06).
1.1 Related literature
A two-class linear classifier seeks to separate all Red-Blue points in the input, thus the focus is on minimizing the mean-square error, which is a tractable problem. In our problem, we assume that either a low error linear classifier does not exist, or even if it exists, it is not supported over a small set of co-ordinates. Thus, the challenge in our problem is to identify the correct subset. It resembles a version of the MAX-CUT problem GW (95) – we want to identify a linear classifier that cuts as many Red-Blue edges as possible, and ensures that the side with Red points has relatively few Blue points. Moreover, we want most of the coefficients of classifier to be close to zero. This class of problems are not as tractable – most versions of the problem are computationally hard (cf. balanced cut, sparsest cut ARV (09) and so on). However, it turns out that efficient (similar in efficiency to computing an optimal linear classifier) approximate algorithms can and do exist. Moreover, the approximation guarantee of some such algorithms is pretty close to $1$ (for e.g. GW (95)). In Subsection 2.1, we show that formulating our problem as a 00-$1$ optimization problem (similar to MAX-CUT), and using a SDP based rounding algorithm, allows us to achieve similar efficiency as a linear classifier, and also obtain desired sensitivity vs specificity trade-offs.
Finally, while algorithms for cut problems, like balanced cut, have been used for clustering BLUvB (13), they are not specifically designed to account for inputs embedded in $\mathbb{R}^{d}$. On the other hand, one can think of the identified subcohort as a single homogeneous cluster, and potentially use a standard clustering algorithm operating on inputs in $\mathbb{R}^{d}$ (like Lloyd’s algorithm or other $k$-means algorithms). However, some key difficulties will remain. First, such algorithms compute non-linear separators, which can lead to further loss of explainability. It is unlikely that one can control the solution sparsity [^2] This would be the number of genes used in the test in our application. for such non-linear separators. For example, in the case of Lloyd’s algorithm, the subset of points for the identified subcohort would be bounded by the corresponding Voronoi polytope, which may have an exponential number of facets in the dimension $d$ – leading to a non-linear solution that’s also unlikely to be sparse. Secondly, $k$-means clustering algorithms can be computationally intractable as well – versions of $k$-means that achieve approximation guarantees close to $1+\epsilon$, have time complexity super-exponential in $\frac{1}{\epsilon}$ (see for example KSS (04)). Thirdly, it is unclear how to set the parameters to control the sensitivity and specificity when using a cookie-cutter clustering algorithm for $k$-means. While the size of clusters (proportional to the sensitivity) could be controlled using constraints, it is unclear how to optimize for specificity explicitly in known constrained $k$-means implementations. Finally, while explainable clustering has been studied for $k$-means type algorithms before (for e.g. MDRF (20) and GPSY (23)), such algorithms are restricted to axis-aligned thresholding cuts, as opposed to general but sparse cuts (as in this paper). Thus their techniques and results are significantly different from the ones outlined in this paper.
2 Overview of our results
We want to solve the following high-level problem. Suppose one is given a set of patients which have the same disease but different disease states. For example, in case of cancer patients, some of them may have metastasis but others may not. We want to identify a small set of genes or proteins (or some other markers) that allows us to test and select patients that are in the same state, i.e., we want to design tests that identify predominantly homogeneous subcohorts (subsets) of patients. For example, if roughly 10% of the patient population has metastasis, then we want to design an algorithm that uses a small number of markers to identify a large enough subcohort, say 2% of the overall patient population, of which 70% have metastasis.
Clearly, there are trade-offs involved. If we want the identified subcohort of patients to be a larger fraction of the overall patient population, i.e., if we want the sensitivity to be large, then the probability that most patients in the subcohort will have metastasis, will be smaller, i.e., the specificity will be smaller (equivalently the subcohort will be more inhomogeneous). Furthermore, if we want the set of identifying markers to be small, i.e., if we want the designed test to be sparse, then the specificity is going to be lower (assuming sensitivity is kept constant). On the other hand, if we want to have sparse tests and high specificity in our subcohort, then the sensitivity (measured by subcohort size) will be reduced.
In this paper, we formalize and solve the subcohort identification problem using Algorithm 1,which is a generalization of the Goemans-Williamson semidefinite programming (SDP) algorithm for solving MAX-CUT GW (95). Our theoretical contribution can be summarized as follows:
  • We prove that the ratio of the size of subcohorts that are identified by Algorithm 1 vs. the optimal subcohorts is lower bounded by an absolute constant $\alpha\simeq 0.82$ (see Theorem 2.2 and Subsection 2.1 for details).
The theoretical result should be independently interesting, as it generalizes the Goemans-Williamson calculations. While SDP rounding algorithms for similar problems have been analyzed before (for e.g. FG (95); LLZ (02) and BHPZ (20)), we are not aware of any similar analysis for the problem in our paper. [^3] In a nutshell, the crux of the Goemans-Williamson guarantee relies on computing the maximum ratio of the chord length to the minor arc length in a unit circle, while in our case it hinges on bounding the ratio of the ”perimeter” of a triangle to the perimeter of the circumcircle, the latter is assumed to be of unit radius. It is surprising that a Goemans-Williamson SDP rounding works here, as our objective function is not sign definite. Usually more complicated rounding techniques, like AN (06) are required in such cases.
Furthermore, we empirically study the trade-offs mentioned above, for Algorithm 1, in the context of identifying subcohorts of metastatic breast cancer using the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19). [^4] The dataset contains RNA microarray data of approximately 2,000 breast cancer patients, and methylation data for approximately 13K genes in about 1,500 breast cancer patients. About a thousand patients had one or more positive lymph nodes upon biopsy (metastasis), while the remaining did not (non-metastasis).
For the empirical part, our goal was to identify subcohorts of predominantly metastatic breast cancer cases using tests based on expression of nuclear receptors, i.e., we want the ratio of the metastatic cases to non-metastatic cases to be high for the identified subcohort. We denote this ratio as the specificity of our test, the percentage of metastatic cases in the identified subcohort is denoted as the sensitivity of the test. The number of genes used is roughly proportional to the sparsity parameter of Algorithm 1. Our first empirical result relates the specificity, sensitivity and the number of genes used, as follows:
  • We show that specificity and sensitivity both increase with the number of genes used in the test (see Figure 1(a)). In particular, we obtained subcohorts of size $\sim 20$% of the subsampled patient population, with about 57% specificity, when we use roughly 15 out of 48 nuclear receptors in the designed test.
  • Moreover, if we define our subcohort as an intersection of two subcohorts, from two different runs of Algorithm 1 (over two subsamples of the dataset), then we can increase the specificity by sacrificing on sensitivity (see Figure 1(b)). In particular, we obtained subcohorts of size $\sim 7$% of the subsampled patient population, with about 67% specificity, when we use roughly 45 out of 48 nuclear receptors in the test.
Our second empirical result, is an application of our algorithm to identify statistically significant associations between sets of genes. In this case, we identify statistically significant associations between epigenetic data and downstream nuclear receptor expressions. For example, in the context of the breast cancer dataset, we use subcohorts identified using Algorithm 1 to associate statistically significant changes in methylation of certain known breast cancer related genes to statistically significant changes in expression of certain nuclear receptors (see Figure 3). Identifying statistically significant associations between changes in methylation percentages and changes in nuclear receptor expressions is an interesting problem since methylation of certain genes is known to be a potential cause of progression of breast cancers SK (15), while nuclear receptors are often used as targets of many drugs BSW+ (13). Therefore, we are attempting to link the likely cause of the pathology to a likely treatment target, specific to a subcohort of patients.
The question arises: why is identifying such associations difficult or novel, after all, it involves only correlation computation? A naive average over the entire cohort of patient population, where one indiscriminately averages over non-homogeneous data, may make confidence intervals larger than the separation between the means. Thus it becomes difficult to say whether any identified associations are statistically significant. This is exactly what happens in Figure 2, where we compute the average expression of nuclear receptors and methylation percentages in metastatic and non-metastatic breast cancer patients, over the entire cohort of about 2,000 patients. Based on the figure, one can not differentiate between metastatic vs non-metastatic breast cancer using correlations and simple z-scores of any of the 90 odd genes plotted – the confidence intervals are not narrow enough to separate the means. Similarly, if one computes averages over arbitrary small subsets of the population, across subcohorts, then the confidence intervals are likely to be too large as well. Therefore, we need to average over "good" (homogeneous) subcohorts, i.e., subcohorts where averaging likely leads to narrower confidence intervals but not closely clustered gene expressions (see Subsection 2.2.2 for details). Identifying subcohorts based on a small set of markers should implicitly help ensure homogeneity within the subcohort, and narrow confidence intervals. Using subcohorts identified by Algorithm 1, we obtain the following result for the breast cancer dataset:
  • We used Algorithm 1 to generate several thousands of subcohorts by subsampling the input (one subcohort for each subsampled input). One identified subcohort, together with the resulting statistically significant association between methylation percentages of known breast cancer related genes and nuclear receptor expressions, is shown in Figure 3. For the subcohort in Figure 3, the significant over-expression of LXRB receptor gene suggests an initial hypothesis: patients in the subcohort may be good candidates for LXR-inverse agonists, if supported by clinical evidence.
    Here, it is worthwhile to note that both LXR-inverse agonists CVS+ (19) and LXR agonists HLRL+ (19) have been proposed to be beneficial for various subsets of breast cancer patients. Our algorithm, and its output test, may offer a principled way to identify genomic pathways underlying the disease. Moreover, it may also help identify patients, who could benefit from use of LXR-inverse agonists or LXR agonists (or neither).
Furthermore, in actual practice, the physician may expect to be given genetic sequencing data from a single patient, so it is natural to ask: Is there a way to decide if a given sequenced sample belongs to some "good" subcohort? If it does, the physician may be able to use the computed statistically significant associations between the genes in that subcohort to arrive at an appropriate treatment. So the answer to the question above is likely to be affirmative, as long as one can build a large library that matches a given genetic sequence to that from a precomputed homogeneous subcohort. It’s worth noting here that there are potentially exponentially many subcohorts that one may need to generate, and then identify good subchohorts from amongst them, in order to build a library of homogeneous subcohorts. Thus, one may need to train a neural net to prioritize identification of subcohorts that are more likely to lead to statistically significant associations. We found that the problem is very amenable to learning using neural nets, and for the breast cancer dataset, even a single hidden layer neural net has reasonable accuracy (see Figure 4 and Subsection 2.2.3 for details).
We begin with formal algorithmic description and theoretical details in Subsection 2.1, which is followed by empirical results on the breast cancer dataset in Subsection 2.2. The biologically minded reader may skip Subsection 2.1, while the mathematically minded reader may skip Subsection 2.2, upon first reading.
2.1 Underlying theoretical problem
We begin with an abstract description of our problem. We are given a complete bipartite graph $G\equiv(U,V)$, where the vertices are embedded in a high dimensional space $\mathbb{R}^{d}$. Our goal is to compute a hyperplane $(h,\theta)$, i.e., $\langle h,x\rangle\leq-\theta$, [^5] Here $x\in\mathbb{R}^{d}$ is the input specified embedding of a given vertex in $G$ into $\mathbb{R}^{d}$. with few non-zero coefficients in $h$, that separates a subset $S\equiv(S_{U},S_{V})\subset U\cup V$ of vertices from the rest of the vertices in $G$. In addition, we want $S_{U}$ to be much larger than $S_{V}$, and $S$ to be large.
It is easy to see that the computed subset $S$ satisfying the above requirements is precisely the subcohort meeting the requirements from the previous discussion.
Without loss of generality, we may assume that the input also specifies a vertex $u_{0}\in S_{U}$, since one can always iterate over all vertices in $U$. Note that, $S$ is large and $|S_{U}|\gg|S_{V}|$, together imply that the cut $(S_{U},V\setminus S)$ is large. Therefore, one may formulate the above as a MAX-CUT type integer program, where one associates a ±$\pm 1$ variable $z_{u}$ with each of the $n$ vertices in $G$, and wants to maximize the objective:
$\max_{h\in\mathbb{R}^{d},\atop{\vec{z}\in\{\pm 1\}^{n}}}\sum_{u\in U,\atop{u^{\prime}\in V}}\frac{(z_{u}-z_{u^{\prime}})^{2}+(z_{u_{0}}-z_{u^{\prime}})^{2}-(z_{u}-z_{u_{0}})^{2}}{3}.$ (1)
Maximizing the size of such a cut is equivalent to maximizing the size of $S$, while making $|S_{U}|$ larger than $|S_{V}|$. Without any further constraints, the trivial solution $S_{U}\equiv U,S_{V}\equiv V$ is optimum. However, the separation and sparsity constraints make the optimum non-trivial. These constraints may be enumerated as follows.
  • Separating edges between $U\setminus\{u_{0}\}$ and $V$: These edges should only contribute to the determination of the separating hyperplane if $u\in U\setminus\{u_{0}\}$ is not cut, i.e., $u$ and $u_{0}$ are on the same side ($u$ is in the subcohort). Hence, for every such $u\in U\setminus\{u_{0}\}$, we add the constraint:
    $\langle h,x_{u}\rangle\leq-\theta z_{u}z_{u_{0}}$ (2)
  • Separating edges between $u_{0}$ and $V$: These edges should only contribute to the determination of the separating hyperplane if $u^{\prime}\in V$ is cut, i.e., $u^{\prime}$ and $u_{0}$ are on different sides ($u^{\prime}$ is not in the subcohort). Hence, for every such $u^{\prime}\in V$, we add the constraint:
    $\langle h,x_{u}\rangle\geq\theta(1-z_{u^{\prime}}z_{u_{0}})$ (3)
  • Sparsity: We want the designed test to use few genes, i.e., we want its coefficients to be sparse. Hence we bound the $\ell_{1}$ norm of its coefficients (see for example FR (13)) by an input parameter $s$, which is expressed as:
    $\|h\|_{1}\leq s.$ (4)
Remark 2.1
It is worth noting that the formulated problem above is very similar to maximum margin classification using linear support vector machines. Thus, it is possible to generalize it to non-classifiers using the "kernel trick" (see for example BGV (92)). The only exception being the sparsity constraint above, which needs to be adapted for the chosen kernel. It may be possible to generalize the sparsity constraint for low degree polynomial kernels using standard compressed sensing methods (see for example FR (13)). However, such non-linear extensions are not the focus of this work.
One may relax the discrete objective function and the linear constraints on the embedding into a SDP, as specified below. Note that the SDP relaxation can be efficiently solved using a standard solver unlike the discrete objective function.
$\max_{h\in\mathbb{R}^{d},\atop{\forall i:\ \|v_{i}\|_{2}=1}}$ $\sum_{u\in U,\atop{u^{\prime}\in V}}\frac{2-2\langle v_{u},v_{u^{\prime}}\rangle-2\langle v_{u_{0}},v_{u^{\prime}}\rangle+2\langle v_{u},v_{u_{0}}\rangle}{3},$ (5)
$s.t.$ $\forall u\in U:$ $\ \langle h,x_{u}\rangle\leq-\theta\langle v_{u},v_{u_{0}}\rangle$ (6)
$\forall u^{\prime}\in V:$ $\ \langle h,x_{u^{\prime}}\rangle\geq\theta(1-\langle v_{u^{\prime}},v_{u_{0}}\rangle)$ (8)
$\|h\|_{1}\leq s,$
where $x_{u}$ are the embeddings in $\mathbb{R}^{d}$, $s$ is the sparsity parameter, and $\theta$ is the separation threshold parameter, all to be specified with the input. Algorithm 1 essentially maps the vector solution of the SDP to ±$\pm 1$ values, to obtain the subcohort (corresponding to $\{i\in U\cup V:\ z_{i}=1\}$).
Algorithm 1 Identify subcohort
1:Input: Two sets of vertices $(U,V)$ labeled 00 and $1$ respectively, embedded in $\mathbb{R}^{d}$; vertex $u_{0}\in U$, sparsity parameter $s\in(0,\infty)$, and separation threshold $\theta\in\mathbb{R}$.
2:Output: If feasible, compute a hyperplane $h\in\mathbb{R}^{d}$ and a subcohort $S\equiv(S_{U},S_{V})\subset U\cup V$ such that: (1) $u_{0}\in S_{u}$, (2) $\forall u\in S_{U}:\ \langle x_{u},h\rangle\leq\theta$ and $\forall v\in S_{V}:\ \langle x_{v},h\rangle\geq 0$, (3) $\|h\|_{1}\leq s$, and (4) $|S_{U}|\gg|S_{V}|$. Note: Here $x_{u}$ and $x_{v}$ are the $\mathbb{R}^{d}$ embeddings corresponding to vertices $u$ and $v$ (from the input) and $\|h\|_{1}$ denotes the $\ell_{1}$ norm of the coefficients of $h$.
3:
4:Solve the SDP relaxation (Equation 5) to obtain the solution vectors $\{v_{i}\}_{n}$.
5:If SDP infeasible, return infeasible.
6:$\forall i\in\{1,...,n\}:\ z_{i}\leftarrow\mathrm{sign}(\langle v_{i},\gamma_{i}\rangle)$, where $\gamma_{i}\sim\mathcal{N}(0,1)$.
7:return $\vec{z},h$
In order to show that the computed SDP solution above is not too far from the optimum, we prove the following approximation guarantee (see the supplement for details).
Theorem 2.2
Let $S^{*}=(S^{*}_{U},S^{*}_{V})$ denote the optimal subcohort, then Algorithm 1 computes a subcohort $S=(S_{U},S_{V})$, such that $|S_{U}|\geq\alpha|S^{*}_{U}|$, where the constant $\alpha\simeq 0.82$, when $|S_{U}^{*}|\ll|U|$.
Note that the optimum hyperplane separator $h$ remains unchanged after rounding though the input specified threshold $\theta$ is perturbed during rounding. However, we may bound the perturbation amount by using Grothendieck’s identity, i.e.,
$\mathbb{E}\left[\mathrm{sign}(z_{u})\mathrm{sign}(z_{u_{0}})\right]=\frac{2}{\pi}\arcsin\left(\langle v_{u},v_{u_{0}}\rangle\right).$ (9)
as follows.
Theorem 2.3
The rounded solution from Algorithm 1 obeys the (slightly weaker) separation constraints:
$\langle h,x_{u}\rangle$$\leq$ $-\frac{2\theta}{\pi}z_{u}z_{u_{0}}$ (10)
$\langle h,x_{u}\rangle$$\geq$ $\theta\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ (11)
as opposed to the corresponding constraints in Equations 2 and 3.
2.2 Application to genomics
In this subsection, we use Algorithm 1 to generate tests [^6] A test is just the output hyperplane of Algorithm 1 – a weighted average of gene expression levels – that evaluates to a score less than the given threshold $\theta$ for metastatic cases, and above $\theta$ for non-metastatic cases. that identifies subcohorts with predominantly metastatic cases in a breast cancer dataset. Our empirical results may be divided into two parts:
  • In Subsection 2.2.1, we examine the quality of the identified subcohorts. A random test can be expected to generate subcohorts with less than 50% metastasis cases (as the number of non-metastasis patients is higher in the dataset). In Figure 1(a), we show that the linear tests generated by Algorithm 1 identify subcohorts where greater than 50% of the cases are metastatic (thus outperforming random). Moreover, if instead of using one hyperplane to identify a subcohort, we use two hyperplanes to identify a subcohort then the identified subcohorts have 60-70% of their cases as metastatic (see Figure 1(b)).
  • In Subsection 2.2.2, we use the identified subcohorts to find statistically significant associations between epigenetic data and gene expressions. In SK (15), the authors identify about 40 genes to be of interest in breast cancers. Many of the genes are tumor suppressors, and methylation of the nucleotides in those genes leads to reduced expression and thus potentially cancer progression. Moreover, in BSW+ (13); OALH (06), the authors discuss the role of 50 nuclear receptors as drug targets. Therefore, it is natural to look for associations between significant changes in nuclear receptor expressions and significant changes in methylation levels of breast cancer related genes (over population baseline levels) – the latter can guide treatment drug targets from the former. We show that, even though at the entire dataset level such associations do not exist, but at the subcohort level they do exist (see Figure 3).
Dataset: We use a subset of the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19) [^7] The dataset is available at: https://www.cbioportal.org/study/summary?id=brca_metabric.that contains RNA microarray data from biopsy samples of approximately 1904 (911 with positive lymph nodes and 993 without positive lymph nodes) breast cancer patients, and methylation data for approximately 13K genes in about 1406 (669 with positive lymph nodes and 737 without positive lymph nodes). In this paper, we (crudely) denote patients with one or more positive lymph nodes upon biopsy as metastasis cases, while the rest as non-metastasis cases.
2.2.1 Predicting subcohorts in breast cancer metastasis
In order to generate Figure 1 (the main result of this subsection), Algorithm 1 takes as input the nuclear receptor expressions of a subset of 50 metastatic and 50 non-metastatic cases from the full dataset, the sparsity parameter $s$, a threshold $\theta$, and an arbitrary metastasis patient that is assumed to belong to the subcohort to be identified. The algorithm outputs a test $h$, where $h$ is just a tuple of weights for the nuclear receptors. On a separate hold-out set of a 1000 patients (500 with metastasis and 500 without), the test simply calculates the weighted average value of the given nuclear receptor expression levels for each patient in the hold-out, and if this value is below the threshold then the patient is added to the subcohort, otherwise not.
Recall that, the sensitivity of Algorithm 1 is the percentage of metastatic cases (out of the hold-out set of 500) that are correctly classified to be in the subcohort, while the specificity is the percentage of metastatic cases in the identified subcohort. The number of genes used by the test is the number of coefficients in the tuple $h$ whose magnitude is within a factor of 10 of the maximum magnitude coefficient in $h$. Due to well known results from compressed sensing theory, the number of genes used should be roughly proportional to the sparsity parameter input to the algorithm.
One expects that as the number of genes increases, so does the specificity and the sensitivity. That is indeed the case in Figure 1. Moreover, one may increase specificity at the cost of sensitivity – simply run Algorithm 1 twice, on two different subsampled sets of patients, and obtain two tests $h_{1},h_{2}$. The new subcohort now consists of the intersection of the previously identified subchorts, i.e., both tests $h_{1},h_{2}$ have to place patients in the subcohort for the patient to be in the new subcohort. Clearly, specificity should increase for the new conservative subcohort compared to the initial subcohorts and sensitivity should decrease. Indeed that is the case – specificity increases to 70% while sensitivity decreases by a factor of 3 (see Figure 1(b)). One may even combine three or more tests in this way. Thus one can use Algorithm 1 to increase the specificity of the subcohort at the cost of decreased sensitivity, i.e., decreased size of the identified subcohort.
(a) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using a single test from Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
(b) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using two tests (as opposed to one), generated using two runs of Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
Figure 1: Sensitivity vs Specificity trade-off using Algorithm 1, for the dataset CSC+ (12). Left: Algorithm 1 outputs a single test hyperplane that identifies subcohorts with 57% specificity, for sparsity parameter $s=6$. Right: When we use the intersection of subcohorts using two tests, the specificity increases to about 70%, but identified subcohort size (sensitivity) decreases to a third, with $s=6$. Note: the single test in the left figure uses about 15 nuclear receptors to identify subcohorts, while the two tests in the right figure use about 45 nuclear receptors.
2.2.2 Associating changes in methylation percentages with changes in nuclear receptors
A central question in genomics is to identify which specific epigenetic changes can lead to clinically significant genetic changes downstream (see for example CH (19)). In this subsection, we use our identified subcohorts to associate epigenetic changes in breast cancer related genes to changes in expression level of nuclear receptors (see Figure 3 for an example). The use of subcohorts is critical because when one averages over the entire dataset the confidence intervals become large (see for example Figure 2), and thus one can not make claims with any degree of statistical confidence.
(a) Expressions of nuclear receptors in metastasis and non-metastasis.
(b) Percentage methylations of breast cancer related genes in metastasis and non-metastasis.
Figure 2: Gene expression and methylation percentages for metastasis and non-metastasis patients computed over the full dataset. Observe that averaging over the entire dataset, consisting of non-homogeneous patient population, leads to large confidence intervals. The latter make it impossible to distinguish between metastasis and non-metastasis using z-scores based on the nuclear receptors and breast cancer related genes used above.
Recall that, methylation of certain tumor suppressor genes can potentially lead to breast cancer progression (see SK (15)), and nuclear receptors are a target for a large number of drugs. Therefore, if we compute the mean expression levels of the two sets of genes over the entire patient population in the dataset, and find a tumor suppressor gene that is significantly more methylated and a nuclear receptor that is significantly over-expressed; then we can test whether a drug targeting that over-expressed nuclear receptor is effective and use it in new patients that show the same pattern of methylation and over-expression. Unfortunately, the situation is not so simple – there are no such sets of tumor suppressors and nuclear receptors where significant change in one is associated with significant change in the other. Figure 2 shows that the confidence intervals are just too large, when we average over the entire dataset (because of inhomogeneity in the disease). [^8] Note: Throughout this paper, the error bars are one standard deviation in length each way.
However, now that we can identify subcohorts of patients, where the identified subcohorts are described by a small set of genes; it is possible that the disease state amongst the patients in the subcohort is homogeneous enough to allow for narrow confidence intervals. That would allow us to identify any significant correlation between changes in tumor suppressor methylation and changes in nuclear receptor expression, amongst patients restricted to a subcohort. Indeed that is the case!
Using Algorithm 1, and thousands of random subsamples of inputs (of 100 patients each), from the original dataset (of size 2000 patients); we generated thousands of patient subcohorts, where each subcohort had a test associated with it. The test could be either based on methylation percentage levels or nuclear receptor expression levels. Next, for every pair of subcohorts, if two subcohorts have more than a certain percentage of patients in common (we used a 10% due to limited dataset size) i.e., they look similar as far as the constituents, then we compute the mean expression and confidence interval for each gene using the common patients. If the subcohorts are made of patients in similar disease states then the confidence intervals should be narrow. And, indeed they are! In Figure 3 we compare expression levels from one such subcohort. Patients in the subcohort are identified by the test, where the coordinate weights (coefficients) for the test hyperplane are given in Figure 3(c,d). This subcohort shows significant methylation increase in BRCA2 and MSH6. Moreover, nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed when compared to the average non-metastatic patient in the dataset. Therefore, LXR-inverse agonists (see for example CVS+ (19)) may prove to be a useful therapeutic target for this subcohort of patients, if indicated by further clinical evidence.
(a) The nuclear receptors in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values. In particular, for patients in this subcohort, NR1H2 is over-expressed. Statistically insignificant bars are not shown.
(b) The methylation percentages in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values. In particular, for patients in this subcohort, BRCA2 and MSH6 were methylated more, both are related to DNA repair, and errors can lead to breast cancer progression SBS+ (19). Statistically insignificant bars are not shown.
(c) The coefficient weights of the genes in the test output by Algorithm 1 to identify the subcohort above.
(d) The coefficient weights of the genes in the test output by Algorithm 1 to identify the subcohort above.
Figure 3: Associating changes in methylation percentages with changes in nuclear receptors using Algorithm 1. Compared to averaging over the entire population (see Figure 2), averaging over patients in a subcohort can lead to statistically significant differences between metastasis and non-metastasis breast cancers. In particular, in the above subcohort of patients, BRCA2 and MSH6 have significantly higher methylation percentages than the entire cohort of the patients in METABRIC dataset; and at the same time they have a significantly higher expression of NR1H2 receptors responsible for lipid metabolism.
2.2.3 Identifying "good" subcohorts
Recall that, searching over every pair of potential subsampled inputs and subcohorts is a computationally intractable problem, since there are an exponential number of such pairs. For the small dataset with 2000 patients, and subsamples of size 1000 patients, we may need to search through $2^{1000}$ runs of Algorithm 1! However, we can train a neural net to predict when we will have subcohorts of patients that are likely to lead to statistically significant associations, i.e., the pairs of subcohorts of patients will have enough common patients to lead to narrow confidence intervals. We can then prioritize our confidence interval calculations for such subcohorts. We implemented this prioritization using a simple single hidden layer neural net. We plot the quality of the neural net model, i.e., the errors in Figure 4 in the supplement. Even with a single hidden layer neural network, we see reasonable error margins. Thus prioritization is an easy problem to learn.
3 Conclusion and limitations
In conclusion, we have demonstrated a method to design tests that identify homogeneous subcohorts of patients. This was done by a non-trivial extension of the Goemans-Williamson algorithm GW (95), which may be independently interesting. Furthermore, such subcohorts may be used to design trials that help identify new treatments. In this paper, we have used the identified subcohorts to link epigenetic changes to changes in nuclear receptor expression levels. However, we do not have the resources to verify our predicted associations and any clinical predictions in an actual experimental setting and we hope this work inspires physicians to pursue that.
4 Acknowledgments
The author is grateful to Dr Siddhartha Mukherjee, and members of the Langelier lab (Dr. Natasha Spottiswoode, Dr Van Phan, and Dr Chaz Langlier) for mentioning the underlying clinical problem to him in two different patient-care settings – in patients with AML and sepsis, respectively. The author is also grateful to Dr Jacob Glass for pointing out the public datasets at the cbio portal, and to Professors Subhash Khot, Anupam Gupta and Cliff Stein for helpful comments.

### v2

@@
 Abstract
-We design an efficient algorithm that outputs a linear classifier for identifying homogeneous subsets (equivalently subcohorts) from large inhomogeneous datasets.
-Our theoretical contribution is a rounding technique, similar to that of Goemans and Williamson (1994), that approximates the optimal solution of the underlying optimization problem within a factor of $0.82$.
-As an application, we use our algorithm to design a simple test that can identify homogeneous subcohorts of patients, that are mainly comprised of metastatic cases, from the RNA microarray dataset for breast cancer by Curtis et al.
+We design an efficient algorithm that outputs tests for identifying predominantly homogeneous subcohorts of patients from large in-homogeneous datasets.
+Our theoretical contribution is a rounding technique, similar to that of Goemans and Wiliamson (1995), that approximates the optimal solution within a factor of $0.82$.
+As an application, we use our algorithm to trade-off sensitivity for specificity to systematically identify clinically interesting homogeneous subcohorts of patients in the RNA microarray dataset for breast cancer from Curtis et al.
 (2012).
-Furthermore, we also use the test output by the algorithm to systematically identify subcohorts of patients in which statistically significant changes in methylation levels of tumor suppressor genes co-occur with statistically significant changes in nuclear receptor expression.
-Identifying such homogeneous subcohorts of patients can be useful for the discovery of disease pathways and therapeutics, specific to the subcohort.
+One such clinically interesting subcohort suggests a link between LXR over-expression and BRCA2 and MSH6 methylation levels for patients in that subcohort.
 1 Introduction
 A linear separator in $\mathbb{R}^{d}$ is a basic tool used in statistics and machine learning to partition a set of points into two subsets.
-Such a separator implicitly encodes information about the pairs of points, which it has to separate, in its coefficients.
-Given a set of Red and Blue points to be separated, there may not exist a linear classifier that separates most of the Red and Blue points.
+Given a set of Red and Blue points to be separated, there may not exist a linear separator that separates most of the Red and Blue points.
 Moreover, even if one exists, then it can be expected to have most coefficients to be far from zero, i.e., it’s not sparse, unless there are structural assumptions on the dataset.
-On the other hand, if we want a linear separator that separates a small (but non-trivial) percentage of Red points from Blue points, then we may reasonably expect that a sparse linear separator exists, i.e., one which sets most of its coefficients to zero, and it still has low error.
-This is because the number of Red-Blue pairs separated is smaller, so the amount of information to be encoded in the linear separator is lower.
-Such separators also have the advantage that they are explainable – the role and importance of each feature is clear from its coefficient.
-Therefore, if one assumes the separator is sparse, the focus shifts to identifying a subset of Red points to carve out from the rest, so that: (1) the number of Blue points in the identified subset is relatively small (specificity is high); (2) the relative number of Red points in the subset, compared to the total number of Red points, is large (sensitivity is high); and (3) the number of non-zero coefficients in the separator is small (the separator is sparse).
-Computing such linear separators are the subject of this paper, and we devise a semidefinite programming (SDP) based algorithm (see Algorithm 1) for that purpose.
-Furthermore, we prove theoretical guarantees on the size of the identified subset (see Theorem 2.2).
-One application for such an algorithm is in identifying homogeneous subsets (equivalently subcohorts) of patients in clinical trials.
-Diseases like breast cancer, sepsis and even bacterial infections are fairly inhomogeneous – the same disease may have very different underlying causes (different mutations, different strains of bacteria, and so on), and therefore different treatments.
-No single treatment works for a majority of the patients.
-Therefore, given a drug that works on a small percentage of patients overall, one recourse is to identify a not too small subset for which the drug treatment works with higher likelihood.
-Moreover, one wants the identifying test to rely on a handful of features (few genes, few mutations etc) – to ensure explainability of the test, and also control for sequencing costs in the field.
-This is precisely the abstract problem delineated in the previous paragraph – we are willing to trade sensitivity for specificity.
-Thus, in this paper, we use the algorithm mentioned above (Algorithm 1), on the RNA microarray dataset for breast cancer from CSC+ (12); PCR+ (16); RSS+ (19) [^1] We use this particular dataset because it is large and balanced, i.e., both metastatic and non-metastatic cases are adequately represented.
-Moreover, it has epigenetic data as well as RNA expression data, and linking the two is an interesting open question in biology CH (19).
-to identify subsets of metastatic breast cancer patients that have statistically significant over-expression of one or more nuclear receptors, compared to the full dataset (see Figure 3).
-It’s worth noting here that nuclear receptors are often chosen as drug targets OALH (06).
+On the other hand, if we want a linear separator that separates a small (but non-trivial) percentage of Red points from Blue points, then a sparse linear separator may exist.
+Such sparse separators also have the advantage that they are likely to be interpretable.
+Therefore, the focus shifts to identifying a subset of Red points to carve out from the rest, so that: (1) the number of Blue points in the identified subset is relatively small (specificity is high); (2) the relative number of Red points in the subset, compared to the total number of Red points, is large (sensitivity is high); and (3) the number of non-zero coefficients in the separator is small (the separator is sparse).
+In this paper, we devise a semidefinite programming (SDP) based algorithm (see Algorithm 1) for computing sparse linear separators that identify homogeneous subgroups.
+We prove theoretical guarantees on the specificity and sensitivity of the separator (see Theorems 3.2 and 3.3).
+We apply our algorithm to design a test for identifying subgroups of metastatic breast cancer cases from a large breast cancer dataset metabric; metabric2; metabric3.
+We empirically study the trade-offs between the sensitivity, specificity and sparsity of the tests output by our algorithm in Section 4.1.
+Finally, in Section 4.2, we use our algorithm to identify clinically interesting subgroups of breast cancer patients.
 1.1 Related literature
-A two-class linear classifier seeks to separate all Red-Blue points in the input, thus the focus is on minimizing the mean-square error, which is a tractable problem.
-In our problem, we assume that either a low error linear classifier does not exist, or even if it exists, it is not supported over a small set of co-ordinates.
-Thus, the challenge in our problem is to identify the correct subset.
-It resembles a version of the MAX-CUT problem GW (95) – we want to identify a linear classifier that cuts as many Red-Blue edges as possible, and ensures that the side with Red points has relatively few Blue points.
-Moreover, we want most of the coefficients of classifier to be close to zero.
-This class of problems are not as tractable – most versions of the problem are computationally hard (cf.
-balanced cut, sparsest cut ARV (09) and so on).
-However, it turns out that efficient (similar in efficiency to computing an optimal linear classifier) approximate algorithms can and do exist.
-Moreover, the approximation guarantee of some such algorithms is pretty close to $1$ (for e.g.
-GW (95)).
-In Subsection 2.1, we show that formulating our problem as a 00-$1$ optimization problem (similar to MAX-CUT), and using a SDP based rounding algorithm, allows us to achieve similar efficiency as a linear classifier, and also obtain desired sensitivity vs specificity trade-offs.
-Finally, while algorithms for cut problems, like balanced cut, have been used for clustering BLUvB (13), they are not specifically designed to account for inputs embedded in $\mathbb{R}^{d}$.
-On the other hand, one can think of the identified subcohort as a single homogeneous cluster, and potentially use a standard clustering algorithm operating on inputs in $\mathbb{R}^{d}$ (like Lloyd’s algorithm or other $k$-means algorithms).
-However, some key difficulties will remain.
-First, such algorithms compute non-linear separators, which can lead to further loss of explainability.
-It is unlikely that one can control the solution sparsity [^2] This would be the number of genes used in the test in our application.
-for such non-linear separators.
-For example, in the case of Lloyd’s algorithm, the subset of points for the identified subcohort would be bounded by the corresponding Voronoi polytope, which may have an exponential number of facets in the dimension $d$ – leading to a non-linear solution that’s also unlikely to be sparse.
-Secondly, $k$-means clustering algorithms can be computationally intractable as well – versions of $k$-means that achieve approximation guarantees close to $1+\epsilon$, have time complexity super-exponential in $\frac{1}{\epsilon}$ (see for example KSS (04)).
-Thirdly, it is unclear how to set the parameters to control the sensitivity and specificity when using a cookie-cutter clustering algorithm for $k$-means.
-While the size of clusters (proportional to the sensitivity) could be controlled using constraints, it is unclear how to optimize for specificity explicitly in known constrained $k$-means implementations.
-Finally, while explainable clustering has been studied for $k$-means type algorithms before (for e.g.
-MDRF (20) and GPSY (23)), such algorithms are restricted to axis-aligned thresholding cuts, as opposed to general but sparse cuts (as in this paper).
-Thus their techniques and results are significantly different from the ones outlined in this paper.
+Perhaps the most well-known subgroup identification algorithm is the Patient Rule Induction Method (PRIM) ff and its variants rlbk; prim-cart.
+These are local search algorithms with heuristics designed to gradually trade-off sensitivity for specificity.
+At it’s crux, the better-known PRIM algorithm relies on identifying one or more hyper-rectangles, where points in the interior of a hyper-rectangle constitute the identified subgroup.
+While such bump hunting heuristics are hard to analyze theoretically, explainable clustering algorithms using axis-aligned thresholding cuts have been theoretically analyzed (for example SIELING2008394expl-kmeans1 and expl-kmeans2).
+There are at least two significant differences between algorithms using axis aligned hyper-rectangles (equivalent to thresholding rules on coordinates) and our approach.
+First, such rule-based algorithms are almost always restricted to axis-aligned thresholding cuts, as opposed to sparse cuts (as in this paper); even though the latter are also intuitively interpretable.
+Second, each axis aligned half-space can also be thought of as node in a decision tree, so computing subgroups that are identified by one or more axis-aligned hyper-rectangles corresponds to computing a decision tree, where the underlying objective is chosen to balance sensitivity and specificity.
+It is well-known that computing optimal decision trees for most reasonable objectives is not just a NP-hard problem, but most versions are even hard to approximate within any constant factor (see for e.g.
+gupta1; koch2024superconstant).
+Intuitively, this is because of close connections between decision trees and circuit complexity.
+In contrast, in this paper, we translate subgroup identification with sparse linear cuts to a MAX-CUT like problem (cf.
+gw), which admits constant factor approximation guarantees via convex optimization (an approximation factor of $0.82$ in our case).
+Moreover, there are practical examples where a sparse linear hyperplane, as opposed to a rule based algorithm, is used to identify interesting subcohorts.
+For example, the papers fbtest and meld describe diagnostic scores, essentially sparse regression models, that are widely used in gastroenterology.
 2 Overview of our results
-We want to solve the following high-level problem.
-Suppose one is given a set of patients which have the same disease but different disease states.
+Motivation: Suppose we are given a set of patients in different disease states.
 For example, in case of cancer patients, some of them may have metastasis but others may not.
-We want to identify a small set of genes or proteins (or some other markers) that allows us to test and select patients that are in the same state, i.e., we want to design tests that identify predominantly homogeneous subcohorts (subsets) of patients.
-For example, if roughly 10% of the patient population has metastasis, then we want to design an algorithm that uses a small number of markers to identify a large enough subcohort, say 2% of the overall patient population, of which 70% have metastasis.
+We want to identify a small set of genes or proteins (or some other markers) that allows us to test and select patients that are in the same state, i.e., we want to design tests that identify predominantly homogeneous subcohorts (subsets) of breast cancer patients with metastasis.
+For example, if roughly 10% of the cancer patients have metastasis, then we want to design an algorithm that uses a small number of gene expressions to identify a large enough subcohort, say 2% of the overall patient population, of which 70% have metastasis.
 Clearly, there are trade-offs involved.
@@
 Furthermore, if we want the set of identifying markers to be small, i.e., if we want the designed test to be sparse, then the specificity is going to be lower (assuming sensitivity is kept constant).
-On the other hand, if we want to have sparse tests and high specificity in our subcohort, then the sensitivity (measured by subcohort size) will be reduced.
-In this paper, we formalize and solve the subcohort identification problem using Algorithm 1,which is a generalization of the Goemans-Williamson semidefinite programming (SDP) algorithm for solving MAX-CUT GW (95).
+On the other hand, if we want to have sparse tests and high specificity in our subcohort, then the sensitivity (which corresponds to subcohort size) will be reduced.
+Contribution: In this paper, we formalize and solve the subcohort identification problem using Algorithm 1, which is a generalization of the Goemans-Williamson semidefinite programming (SDP) algorithm for solving MAX-CUT gw.
 Our theoretical contribution can be summarized as follows:
   • We prove that the ratio of the size of subcohorts that are identified by Algorithm 1 vs.
-the optimal subcohorts is lower bounded by an absolute constant $\alpha\simeq 0.82$ (see Theorem 2.2 and Subsection 2.1 for details).
+the optimal subcohorts is lower bounded by an absolute constant $\alpha\simeq 0.82$ (see Section 3 for details).
+This implies bounds on the sensitivity and specificity as well (Theorems 3.2 and 3.3).
 The theoretical result should be independently interesting, as it generalizes the Goemans-Williamson calculations.
-While SDP rounding algorithms for similar problems have been analyzed before (for e.g.
-FG (95); LLZ (02) and BHPZ (20)), we are not aware of any similar analysis for the problem in our paper.
+While similar SDP rounding algorithms for similar problems have been analyzed before (for e.g.
+feige1; rpr2 and potechin), we are not aware of any similar analysis for the problem in our paper.
 [^1] In a nutshell, the crux of the Goemans-Williamson guarantee relies on computing the maximum ratio of the chord length to the minor arc length in a unit circle, while in our case it hinges on bounding the ratio of the ”perimeter” of a triangle to the perimeter of the circumcircle, the latter is assumed to be of unit radius.
 It is surprising that a Goemans-Williamson SDP rounding works here, as our objective function is not sign definite.
-Usually more complicated rounding techniques, like AN (06) are required in such cases.
-Furthermore, we empirically study the trade-offs mentioned above, for Algorithm 1, in the context of identifying subcohorts of metastatic breast cancer using the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19).
-[^4] The dataset contains RNA microarray data of approximately 2,000 breast cancer patients, and methylation data for approximately 13K genes in about 1,500 breast cancer patients.
+Usually more complicated rounding techniques, like alon-naor are required in such cases.
+Furthermore, we empirically study the trade-offs mentioned above, for Algorithm 1, in the context of identifying subcohorts of metastatic breast cancer using the METABRIC dataset metabric; metabric2; metabric3.222The dataset contains RNA microarray data of approximately 2,000 breast cancer patients, and methylation data for approximately 13K genes in about 1,500 breast cancer patients.
 About a thousand patients had one or more positive lymph nodes upon biopsy (metastasis), while the remaining did not (non-metastasis).
-For the empirical part, our goal was to identify subcohorts of predominantly metastatic breast cancer cases using tests based on expression of nuclear receptors, i.e., we want the ratio of the metastatic cases to non-metastatic cases to be high for the identified subcohort.
-We denote this ratio as the specificity of our test, the percentage of metastatic cases in the identified subcohort is denoted as the sensitivity of the test.
+For the empirical part, our first result identifies subcohorts of predominantly metastatic breast cancer cases using sparse linear tests based on expression of nuclear receptors, i.e., we want the ratio of the metastatic cases to all breast cancer cases to be high for the identified subcohort.
+We denote this ratio as the specificity of our test.
+The number of metastatic cases in the identified subcohort as a percentage of all breast cancer cases is denoted as the sensitivity of the test.
 The number of genes used is roughly proportional to the sparsity parameter of Algorithm 1.
 Our first empirical result relates the specificity, sensitivity and the number of genes used, as follows:
-  • We show that specificity and sensitivity both increase with the number of genes used in the test (see Figure 1(a)).
+  • We show that specificity and sensitivity both increase with the number of genes used in the test (see Figure 1(a) and Subsection 4.1 for details).
 In particular, we obtained subcohorts of size $\sim 20$% of the subsampled patient population, with about 57% specificity, when we use roughly 15 out of 48 nuclear receptors in the designed test.
-  • Moreover, if we define our subcohort as an intersection of two subcohorts, from two different runs of Algorithm 1 (over two subsamples of the dataset), then we can increase the specificity by sacrificing on sensitivity (see Figure 1(b)).
+  • Moreover, if we define our subcohort as an intersection of half-spaces from two different runs of Algorithm 1 (over two subsamples of the dataset), then we can increase the specificity by sacrificing on sensitivity (see Figure 1(b)).
 In particular, we obtained subcohorts of size $\sim 7$% of the subsampled patient population, with about 67% specificity, when we use roughly 45 out of 48 nuclear receptors in the test.
-Our second empirical result, is an application of our algorithm to identify statistically significant associations between sets of genes.
-In this case, we identify statistically significant associations between epigenetic data and downstream nuclear receptor expressions.
-For example, in the context of the breast cancer dataset, we use subcohorts identified using Algorithm 1 to associate statistically significant changes in methylation of certain known breast cancer related genes to statistically significant changes in expression of certain nuclear receptors (see Figure 3).
-Identifying statistically significant associations between changes in methylation percentages and changes in nuclear receptor expressions is an interesting problem since methylation of certain genes is known to be a potential cause of progression of breast cancers SK (15), while nuclear receptors are often used as targets of many drugs BSW+ (13).
-Therefore, we are attempting to link the likely cause of the pathology to a likely treatment target, specific to a subcohort of patients.
-The question arises: why is identifying such associations difficult or novel, after all, it involves only correlation computation?
+Our second empirical result, is an application of Algorithm 1 to identify clinically interesting large homogeneous subcohorts of patients.
+One example of a clinically interesting problem is to identify statistically significant associations between changes in methylation percentages and changes in nuclear receptor expressions.
+Methylation of tumor suppressor genes is known to be a potential cause of progression of breast cancers brcgenes, while nuclear receptors are often used as targets of many drugs nrdrugs.
+Therefore, we are attempting to associate the likely cause of the pathology to a likely treatment target, for a specific subcohort of patients.
+The question arises: why is identifying such associations difficult or novel, after all, it involves only standard deviation computation?
 A naive average over the entire cohort of patient population, where one indiscriminately averages over non-homogeneous data, may make confidence intervals larger than the separation between the means.
-Thus it becomes difficult to say whether any identified associations are statistically significant.
+Thus deviations from the mean are no longer statistically significant.
 This is exactly what happens in Figure 2, where we compute the average expression of nuclear receptors and methylation percentages in metastatic and non-metastatic breast cancer patients, over the entire cohort of about 2,000 patients.
-Based on the figure, one can not differentiate between metastatic vs non-metastatic breast cancer using correlations and simple z-scores of any of the 90 odd genes plotted – the confidence intervals are not narrow enough to separate the means.
-Similarly, if one computes averages over arbitrary small subsets of the population, across subcohorts, then the confidence intervals are likely to be too large as well.
-Therefore, we need to average over "good" (homogeneous) subcohorts, i.e., subcohorts where averaging likely leads to narrower confidence intervals but not closely clustered gene expressions (see Subsection 2.2.2 for details).
-Identifying subcohorts based on a small set of markers should implicitly help ensure homogeneity within the subcohort, and narrow confidence intervals.
-Using subcohorts identified by Algorithm 1, we obtain the following result for the breast cancer dataset:
-  • We used Algorithm 1 to generate several thousands of subcohorts by subsampling the input (one subcohort for each subsampled input).
-One identified subcohort, together with the resulting statistically significant association between methylation percentages of known breast cancer related genes and nuclear receptor expressions, is shown in Figure 3.
-For the subcohort in Figure 3, the significant over-expression of LXRB receptor gene suggests an initial hypothesis: patients in the subcohort may be good candidates for LXR-inverse agonists, if supported by clinical evidence.
-    Here, it is worthwhile to note that both LXR-inverse agonists CVS+ (19) and LXR agonists HLRL+ (19) have been proposed to be beneficial for various subsets of breast cancer patients.
+Based on the figure, one can not differentiate between metastatic vs non-metastatic breast cancer using deviations and simple Z-scores of any of the 90 odd genes plotted – the population is inhomogeneous – confidence intervals are too wide to differentiate between the two classes of patients with statistical significance.
+Similarly, if one computes averages over many small subsets of the population, across subcohorts, then the confidence intervals are likely to be too large due to small sample size.
+Therefore, to deduce clinically interesting conclusions, we need to identify and compute the deviations over homogeneous subcohorts.
+In Subsection 4.2, we trade-off sensitivity for specificity using Algorithm 1 to make the identified subcohorts more homogeneous.
+One clinically interesting subcohort is shown in Figure 3.
+For this subcohort, we observe:
+  • There is significant methylation increase in genes BRCA2 and MSH6, and nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed, both relative to the baseline population.
+Therefore, one may hypothesize that methylation of the former leads to downstream over-expression of the latter.
+  • Furthermore, significant over-expression of LXRB receptor gene suggests that patients in the subcohort may be good candidates for LXR-inverse agonists, if supported by clinical evidence.
+    Here, it is worthwhile to note that both LXR-inverse agonists lxr-ant and LXR agonists lxrbeta have been proposed to be beneficial for various subsets of breast cancer patients.
 Our algorithm, and its output test, may offer a principled way to identify genomic pathways underlying the disease.
 Moreover, it may also help identify patients, who could benefit from use of LXR-inverse agonists or LXR agonists (or neither).
-Furthermore, in actual practice, the physician may expect to be given genetic sequencing data from a single patient, so it is natural to ask: Is there a way to decide if a given sequenced sample belongs to some "good" subcohort?
-If it does, the physician may be able to use the computed statistically significant associations between the genes in that subcohort to arrive at an appropriate treatment.
-So the answer to the question above is likely to be affirmative, as long as one can build a large library that matches a given genetic sequence to that from a precomputed homogeneous subcohort.
-It’s worth noting here that there are potentially exponentially many subcohorts that one may need to generate, and then identify good subchohorts from amongst them, in order to build a library of homogeneous subcohorts.
-Thus, one may need to train a neural net to prioritize identification of subcohorts that are more likely to lead to statistically significant associations.
-We found that the problem is very amenable to learning using neural nets, and for the breast cancer dataset, even a single hidden layer neural net has reasonable accuracy (see Figure 4 and Subsection 2.2.3 for details).
-We begin with formal algorithmic description and theoretical details in Subsection 2.1, which is followed by empirical results on the breast cancer dataset in Subsection 2.2.
-The biologically minded reader may skip Subsection 2.1, while the mathematically minded reader may skip Subsection 2.2, upon first reading.
-2.1 Underlying theoretical problem
+Currently, clinically homogeneous subgroups of breast cancer patients are often identified by single genes.
+For example, ER+ patients, i.e., those breast cancer patients with ER nuclear receptor expression above a certain threshold, are treated with a certain protocol and vice versa bc-protocol.
+This is a likely suboptimal determination of patient subcohorts and resulting treatment protocols.
+In contrast, our algorithm presents a systematic approach with a potentially better handle on sensitivity, specificity and sparsity trade-offs.
+Organization: We begin with formal algorithmic description and theoretical details in Section 3, which is followed by empirical results on the breast cancer dataset in Section 4.
+3 Underlying theoretical problem
 We begin with an abstract description of our problem.
@@
 Note that, $S$ is large and $|S_{U}|\gg|S_{V}|$, together imply that the cut $(S_{U},V\setminus S)$ is large.
-Therefore, one may formulate the above as a MAX-CUT type integer program, where one associates a ±$\pm 1$ variable $z_{u}$ with each of the $n$ vertices in $G$, and wants to maximize the objective:
+Therefore, one may formulate the above as a MAX-CUT type integer program, where one associates a $\pm 1$ variable $z_{u}$ with each of the $n$ vertices in $G$, and wants to maximize the objective:
 $\max_{h\in\mathbb{R}^{d},\atop{\vec{z}\in\{\pm 1\}^{n}}}\sum_{u\in U,\atop{u^{\prime}\in V}}\frac{(z_{u}-z_{u^{\prime}})^{2}+(z_{u_{0}}-z_{u^{\prime}})^{2}-(z_{u}-z_{u_{0}})^{2}}{3}.$ (1)
@@
 Hence, for every such $u^{\prime}\in V$, we add the constraint:
-    $\langle h,x_{u}\rangle\geq\theta(1-z_{u^{\prime}}z_{u_{0}})$ (3)
+    $\langle h,x_{u^{\prime}}\rangle\geq\theta(1-z_{u^{\prime}}z_{u_{0}})$ (3)
   • Sparsity: We want the designed test to use few genes, i.e., we want its coefficients to be sparse.
-Hence we bound the $\ell_{1}$ norm of its coefficients (see for example FR (13)) by an input parameter $s$, which is expressed as:
+Hence we bound the $\ell_{1}$ norm of its coefficients (see for example rauhut) by an input parameter $s$, which is expressed as:
     $\|h\|_{1}\leq s.$ (4)
@@
 It is worth noting that the formulated problem above is very similar to maximum margin classification using linear support vector machines.
-Thus, it is possible to generalize it to non-classifiers using the "kernel trick" (see for example BGV (92)).
+Thus, it is possible to generalize it to non-classifiers using the "kernel trick" (see for example vapnik).
 The only exception being the sparsity constraint above, which needs to be adapted for the chosen kernel.
-It may be possible to generalize the sparsity constraint for low degree polynomial kernels using standard compressed sensing methods (see for example FR (13)).
+It may be possible to generalize the sparsity constraint for low degree polynomial kernels using standard compressed sensing methods (see for example rauhut).
 However, such non-linear extensions are not the focus of this work.
@@
 ${s.t.}$ ${\forall u\in U:}$ $\ {\langle h,x_{u}\rangle\leq-\theta\langle v_{u},v_{u_{0}}\rangle}$ (6)
-$\forall u^{\prime}\in V:$ $\ \langle h,x_{u^{\prime}}\rangle\geq\theta(1-\langle v_{u^{\prime}},v_{u_{0}}\rangle)$ (8)
+${\forall u^{\prime}\in V:}$ $\ {\scriptstyle\langle h,x_{u^{\prime}}\rangle\geq\theta(1-\langle v_{u^{\prime}},v_{u_{0}}\rangle)}$ (8)
 ${\|h\|_{1}\leq s,}$
 where $x_{u}$ are the embeddings in $\mathbb{R}^{d}$, $s$ is the sparsity parameter, and $\theta$ is the separation threshold parameter, all to be specified with the input.
-Algorithm 1 essentially maps the vector solution of the SDP to ±$\pm 1$ values, to obtain the subcohort (corresponding to $\{i\in U\cup V:\ z_{i}=1\}$).
+Algorithm 1 essentially maps the vector solution of the SDP to $\pm 1$ values, to obtain the subcohort (corresponding to $\{i\in U\cup V:\ z_{i}=1\}$).
 Algorithm 1 Identify subcohort
-1:Input: Two sets of vertices $(U,V)$ labeled 00 and $1$ respectively, embedded in $\mathbb{R}^{d}$; vertex $u_{0}\in U$, sparsity parameter $s\in(0,\infty)$, and separation threshold $\theta\in\mathbb{R}$.
+1:Input: Two sets of vertices $(U,V)$ labeled 0 and $1$ respectively, embedded in $\mathbb{R}^{d}$; vertex $u_{0}\in U$, sparsity parameter $s\in(0,\infty)$, and separation threshold $\theta\in\mathbb{R}$.
 2:Output: If feasible, compute a hyperplane $h\in\mathbb{R}^{d}$ and a subcohort $S\equiv(S_{U},S_{V})\subset U\cup V$ such that: (1) $u_{0}\in S_{U}$, (2) $\forall u\in S_{U}:\ \langle x_{u},h\rangle\leq\theta$ and $\forall v\in S_{V}:\ \langle x_{v},h\rangle\geq 0$, (3) $\|h\|_{1}\leq s$, and (4) $|S_{U}|\gg|S_{V}|$.
@@
 3:
-4:Solve the SDP relaxation (Equation 5) to obtain the solution vectors $\{v_{i}\}_{n}$.
+4:Solve the SDP relaxation (Equation 5) to obtain the solution vectors $\{v_{i}\}_{i\in[n]}$.
 5:If SDP infeasible, return infeasible.
-6:$\forall i\in\{1,...,n\}:\ z_{i}\leftarrow\mathrm{sign}(\langle v_{i},\gamma_{i}\rangle)$, where $\gamma_{i}\sim\mathcal{N}(0,1)$.
+[…]
 7:return $\vec{z},h$
-In order to show that the computed SDP solution above is not too far from the optimum, we prove the following approximation guarantee (see the supplement for details).
-Theorem 2.2
-Let $S^{*}=(S^{*}_{U},S^{*}_{V})$ denote the optimal subcohort, then Algorithm 1 computes a subcohort $S=(S_{U},S_{V})$, such that $|S_{U}|\geq\alpha|S^{*}_{U}|$, where the constant $\alpha\simeq 0.82$, when $|S_{U}^{*}|\ll|U|$.
-Note that the optimum hyperplane separator $h$ remains unchanged after rounding though the input specified threshold $\theta$ is perturbed during rounding.
+In order to show that the sensitivity and specificity of the test output by the SDP solution above are not too far from the optimum, we prove the following approximation guarantee (see the appendix, at the end, for proofs).
+Theorem 3.2 (Sensitivity)
+Let $S^{*}=(S^{*}_{U},S^{*}_{V})$ denote the optimal subcohort such that $|S_{U}^{*}|\ll|U|$, then Algorithm 1 computes a subcohort $S=(S_{U},S_{V})$, such that $|S_{U}|\geq\alpha|S^{*}_{U}|$, where the constant $\alpha\simeq 0.82$.
+Equivalently, $\frac{|S_{U}|}{|U|+|V|}\geq\alpha\cdot\frac{|S^{*}_{U}|}{|U|+|V|}$, i.e., the sensitivity of the output test is within $0.82$ of the optimum.
+Theorem 3.3 (Specificity)
+For $\kappa\in(0,1)$, let $\kappa\cdot|U||V|$ denote the fraction of edges in the optimal sparse cut separating $U$ from $V$.
+The specificity of the test output by Algorithm 1 is lower bounded by the solution to the following optimization problem:
+$\min_{x,y\in(0,1)}\frac{x\cdot|U|}{(1-y)|V|+x|U|}\qquad\mathrm{s.t.}\quad x\cdot y\geq\tilde{\alpha}\cdot\kappa,$ (9)
+where $\tilde{\alpha}=\frac{\alpha}{2}\simeq 0.41$.
+Note that the solution of the optimization problem in Equation 9 grows as $\Omega(\sqrt{\kappa})$, when $|U|$ is assumed equal to $|V|$ (this can be seen by substituting $y=\frac{\tilde{\alpha}\cdot\kappa}{x}$ in the objective).
+A plot of the solution surface, together with the proof, is given in the appendix (Section 7).
+While the optimum hyperplane separator $h$ remains unchanged after rounding though the input specified threshold $\theta$ is perturbed during rounding.
 However, we may bound the perturbation amount by using Grothendieck’s identity, i.e.,
@@
 $\langle h,x_{u}\rangle$ $\leq$ $-\frac{2\theta}{\pi}z_{u}z_{u_{0}}$ (11)
-$\langle h,x_{u}\rangle$$\geq$ $\theta\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ (11)
+$\langle h,x_{u^{\prime}}\rangle$ $\geq$ $\theta\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ (12)
 as opposed to the corresponding constraints in Equations 2 and 3.
-2.2 Application to genomics
+4 Experiments: Genomics applications
 In this subsection, we use Algorithm 1 to generate tests [^4] A test is just the output hyperplane of Algorithm 1 – a weighted average of gene expression levels – that evaluates to a score less than the given threshold $\theta$ for metastatic cases, and above $\theta$ for non-metastatic cases.
@@
 Our empirical results may be divided into two parts:
-  • In Subsection 2.2.1, we examine the quality of the identified subcohorts.
-A random test can be expected to generate subcohorts with less than 50% metastasis cases (as the number of non-metastasis patients is higher in the dataset).
+  • In Subsection 4.1, we examine the quality of the identified subcohorts.
+A random test, i.e., a random half-space, can be expected to generate subcohorts with less than 50% metastasis cases (as the number of non-metastasis patients is higher in the dataset).
 In Figure 1(a), we show that the linear tests generated by Algorithm 1 identify subcohorts where greater than 50% of the cases are metastatic (thus outperforming random).
 Moreover, if instead of using one hyperplane to identify a subcohort, we use two hyperplanes to identify a subcohort then the identified subcohorts have 60-70% of their cases as metastatic (see Figure 1(b)).
-  • In Subsection 2.2.2, we use the identified subcohorts to find statistically significant associations between epigenetic data and gene expressions.
-In SK (15), the authors identify about 40 genes to be of interest in breast cancers.
-Many of the genes are tumor suppressors, and methylation of the nucleotides in those genes leads to reduced expression and thus potentially cancer progression.
-Moreover, in BSW+ (13); OALH (06), the authors discuss the role of 50 nuclear receptors as drug targets.
+We also compare our algorithm’s specificity (for a given sensitivity) with a version of the well-known PRIM algorithm for subgroup discovery (this comparison is in the Appendix Section 9).
+  • In Subsection 4.2, we use Algorithm 1 to identify a clinically interesting subcohort of breast cancer patients – those that may respond to LXR inverse agonists.
+In the process, we also associate epigenetic changes with changes in nuclear receptor expressions.
+It is known brcgenes, that about $40$ genes act as tumor suppressors, and their methylation leads to reduced expression and potentially cancer progression.
+Moreover, in nrdrugs; nrdrugs2, the authors discuss the role of nuclear receptors as drug targets.
 Therefore, it is natural to look for associations between significant changes in nuclear receptor expressions and significant changes in methylation levels of breast cancer related genes (over population baseline levels) – the latter can guide treatment drug targets from the former.
 We show that, even though at the entire dataset level such associations do not exist, but at the subcohort level they do exist (see Figure 3).
-Dataset: We use a subset of the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19) [^7] The dataset is available at: https://www.cbioportal.org/study/summary?id=brca_metabric.that contains RNA microarray data from biopsy samples of approximately 1904 (911 with positive lymph nodes and 993 without positive lymph nodes) breast cancer patients, and methylation data for approximately 13K genes in about 1406 (669 with positive lymph nodes and 737 without positive lymph nodes).
+Dataset: The results in this section are based on the METABRIC dataset metabric; metabric2; metabric3, which contains RNA microarray data from biopsy samples of approximately 1904 (911 with positive lymph nodes and 993 without positive lymph nodes) breast cancer patients, and methylation data for approximately 13K genes in about 1406 (669 with positive lymph nodes and 737 without positive lymph nodes).
 In this paper, we (crudely) denote patients with one or more positive lymph nodes upon biopsy as metastasis cases, while the rest as non-metastasis cases.
-2.2.1 Predicting subcohorts in breast cancer metastasis
-In order to generate Figure 1 (the main result of this subsection), Algorithm 1 takes as input the nuclear receptor expressions of a subset of 50 metastatic and 50 non-metastatic cases from the full dataset, the sparsity parameter $s$, a threshold $\theta$, and an arbitrary metastasis patient that is assumed to belong to the subcohort to be identified.
+4.1 Predicting breast cancer metastasis
+Recall that, the sensitivity of Algorithm 1 is the percentage of metastatic cases that are correctly classified to be in the subcohort, while the specificity is the percentage of metastatic cases in the identified subcohort.
+The number of genes used by the test is defined as the number of coefficients of $h$ whose magnitude is at least $\frac{1}{10}$ of the maximum magnitude coefficient in $h$.
+Due to well known results from $\ell_{1}$ norm minimization rauhut, the number of genes used should be roughly proportional to the sparsity parameter input to the algorithm (Equation 4).
+Figure 1 shows the sensitivity, specificity and sparsity trade-offs for our subcohort identification algorithm.
+To generate Figure 1, Algorithm 1 uses the nuclear receptor expressions of a subset of 50 metastatic and 50 non-metastatic cases from the full dataset, the sparsity parameter $s$, a threshold $\theta$, and an arbitrary metastasis patient that is assumed to belong to the subcohort to be identified.
 The algorithm outputs a test $h$, where $h$ is just a tuple of weights for the nuclear receptors.
 On a separate hold-out set of a 1000 patients (500 with metastasis and 500 without), the test simply calculates the weighted average value of the given nuclear receptor expression levels for each patient in the hold-out, and if this value is below the threshold then the patient is added to the subcohort, otherwise not.
-Recall that, the sensitivity of Algorithm 1 is the percentage of metastatic cases (out of the hold-out set of 500) that are correctly classified to be in the subcohort, while the specificity is the percentage of metastatic cases in the identified subcohort.
-The number of genes used by the test is the number of coefficients in the tuple $h$ whose magnitude is within a factor of 10 of the maximum magnitude coefficient in $h$.
-Due to well known results from compressed sensing theory, the number of genes used should be roughly proportional to the sparsity parameter input to the algorithm.
-One expects that as the number of genes increases, so does the specificity and the sensitivity.
-That is indeed the case in Figure 1.
-Moreover, one may increase specificity at the cost of sensitivity – simply run Algorithm 1 twice, on two different subsampled sets of patients, and obtain two tests $h_{1},h_{2}$.
-The new subcohort now consists of the intersection of the previously identified subchorts, i.e., both tests $h_{1},h_{2}$ have to place patients in the subcohort for the patient to be in the new subcohort.
-Clearly, specificity should increase for the new conservative subcohort compared to the initial subcohorts and sensitivity should decrease.
+The above exercise is repeated $10$ times for various sparsity values, and sensitivity and specificity measured each time, to arrive at Figure 1.
+Interpretation of Figure 1: First, as the number of genes increases, so do specificity and sensitivity in Figure 1.
+Second, sensitivity can be traded-off for specificity: We run Algorithm 1 twice, on two different subsampled sets of patients, and obtain two half-spaces/tests $h_{1},h_{2}$ that correspond to two subcohorts.
+The new subcohort is defined to be the intersection of the previously identified half-spaces, i.e., both tests $h_{1},h_{2}$ have to place patients in the subcohort for the patient to be in the new subcohort.
+Clearly, specificity should increase for this more conservatively defined subcohort and sensitivity should decrease.
 Indeed that is the case – specificity increases to 70% while sensitivity decreases by a factor of 3 (see Figure 1(b)).
@@
 Thus one can use Algorithm 1 to increase the specificity of the subcohort at the cost of decreased sensitivity, i.e., decreased size of the identified subcohort.
+Comparison: We compare our algorithm with a version of the well-known PRIM algorithm in Appendix (Section 9).
+It is clear from our experiments that the specificity of the MAX-CUT based approach outperforms that of the local search based heuristic when sensitivity is held constant.
+This is despite the fact that the local search heuristic was allowed an order of magnitude more time for computations.
 (a) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using a single test from Algorithm 1.
 Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars.
-The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
-(b) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using two tests (as opposed to one), generated using two runs of Algorithm 1.
+The sparsity parameter (x-axis) is proportional to the number of genes used.
+(b) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using two tests (vs one in L figure), generated using two runs of Algorithm 1.
 Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars.
-The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
-Figure 1: Sensitivity vs Specificity trade-off using Algorithm 1, for the dataset CSC+ (12).
+The sparsity parameter is proportional to the number of genes used.
+Figure 1: Sensitivity vs Specificity trade-off using Algorithm 1, for the dataset metabric.
 Left: Algorithm 1 outputs a single test hyperplane that identifies subcohorts with 57% specificity, for sparsity parameter $s=6$.
@@
 Note: the single test in the left figure uses about 15 nuclear receptors to identify subcohorts, while the two tests in the right figure use about 45 nuclear receptors.
-2.2.2 Associating changes in methylation percentages with changes in nuclear receptors
-A central question in genomics is to identify which specific epigenetic changes can lead to clinically significant genetic changes downstream (see for example CH (19)).
-In this subsection, we use our identified subcohorts to associate epigenetic changes in breast cancer related genes to changes in expression level of nuclear receptors (see Figure 3 for an example).
-The use of subcohorts is critical because when one averages over the entire dataset the confidence intervals become large (see for example Figure 2), and thus one can not make claims with any degree of statistical confidence.
+4.2 Methylation and gene expression
+Algorithm 2 Identify Critical Genes
+1:Input: Dataset $D\subset\mathbb{R}^{d}$, and parameter $\epsilon\in(0,1)$ (the $d$ coordinates are the genes)
+2:Output: A subcohort with critical genes that are significantly differently expressed from the population
+3:
+4:Uniformly sample $S\subset D$ with $|S|\simeq\epsilon\cdot|D|$
+5:Use Algorithm 1 on $S$ to compute a half-space $h$ corresponding to a likely homogeneous subcohort.
+6:Compute means and standard deviation of each of $d$ coordinates in population and subcohort.
+7:If any coordinate has mean which is one standard deviation away from population mean, then return the coordinate (gene), subcohort and $h$ (test)
+8:Else generate a new sample (Step 1) and repeat
+One application of Algorithm 1 is to identify genes which differ significantly between a subcohort and the population.
+Algorithm 2 presents the basic idea.
+We will use a slight variation of it to relate epigenetic changes (nucleotide methylation changes) to clinically significant genetic changes downstream – an important question in genomics (see for example epi1).
+The use of subsets is critical as averaging over the entire dataset leads to large confidence intervals.
 (a) Expressions of nuclear receptors in metastasis and non-metastasis.
@@
 The latter make it impossible to distinguish between metastasis and non-metastasis using z-scores based on the nuclear receptors and breast cancer related genes used above.
-Recall that, methylation of certain tumor suppressor genes can potentially lead to breast cancer progression (see SK (15)), and nuclear receptors are a target for a large number of drugs.
-Therefore, if we compute the mean expression levels of the two sets of genes over the entire patient population in the dataset, and find a tumor suppressor gene that is significantly more methylated and a nuclear receptor that is significantly over-expressed; then we can test whether a drug targeting that over-expressed nuclear receptor is effective and use it in new patients that show the same pattern of methylation and over-expression.
+Background: Methylation of certain tumor suppressor genes can potentially lead to breast cancer progression epi1; brcgenes, and nuclear receptors are a target for a large number of drugs nrdrugs; nrdrugs2.
+Therefore, if we compute the mean methylation and expression levels of critical genes over the entire patient population in the dataset to find a tumor suppressor gene that is significantly more methylated and a nuclear receptor that is significantly over-expressed; then we can test whether a drug targeting that over-expressed nuclear receptor is effective in patients that show a similar pattern of excessive methylation and over-expression.
 Unfortunately, the situation is not so simple – there are no such sets of tumor suppressors and nuclear receptors where significant change in one is associated with significant change in the other.
 Figure 2 shows that the confidence intervals are just too large, when we average over the entire dataset (because of inhomogeneity in the disease).
-[^8] Note: Throughout this paper, the error bars are one standard deviation in length each way.
-However, now that we can identify subcohorts of patients, where the identified subcohorts are described by a small set of genes; it is possible that the disease state amongst the patients in the subcohort is homogeneous enough to allow for narrow confidence intervals.
-That would allow us to identify any significant correlation between changes in tumor suppressor methylation and changes in nuclear receptor expression, amongst patients restricted to a subcohort.
-Indeed that is the case!
-Using Algorithm 1, and thousands of random subsamples of inputs (of 100 patients each), from the original dataset (of size 2000 patients); we generated thousands of patient subcohorts, where each subcohort had a test associated with it.
-The test could be either based on methylation percentage levels or nuclear receptor expression levels.
-Next, for every pair of subcohorts, if two subcohorts have more than a certain percentage of patients in common (we used a 10% due to limited dataset size) i.e., they look similar as far as the constituents, then we compute the mean expression and confidence interval for each gene using the common patients.
-If the subcohorts are made of patients in similar disease states then the confidence intervals should be narrow.
-And, indeed they are!
-In Figure 3 we compare expression levels from one such subcohort.
-Patients in the subcohort are identified by the test, where the coordinate weights (coefficients) for the test hyperplane are given in Figure 3(c,d).
-This subcohort shows significant methylation increase in BRCA2 and MSH6.
+However, it is possible that the disease state amongst the patients in some subcohorts is homogeneous enough to allow for narrow confidence intervals, and thus identify any simultaneous significant deviation in tumor suppressor methylation and nuclear receptor expression, amongst patients restricted to such a subcohort.
+Indeed that is the case below.
+Algorithm 2 for relating epigenetics with nuclear receptors: We use a variant of Algorithm 2, with about $1000$ different random subsamples of inputs (of $100$ patients each), from the original dataset (of size $2000$ patients) to generate patient subcohorts of breast cancer patients (and corresponding tests to identify the subcohorts) in two ways: (1) based on methylation percentage levels, and (2) based on nuclear receptor expression levels.
+Next, for every pair of subcohorts generated (one from methylation levels and another from nuclear receptor expressions), if the two subcohorts have more than a certain percentage of patients in common (we used a threshold of 10% due to limited dataset size) i.e., they share a relatively large subset of the patient population, then they are likely to be homogeneous as two different sets of markers – genetic and epigenetic – are identifying a similar subset of patients.
+Next, we compute the mean methylation and expression together with confidence intervals, for each gene using the common patients.
+Finally, we iterate until we find a patient subcohort with genes that have large deviations from the population means.
+One such subcohort is shown in Figure 3.
+Interpretation of Figure 3: Patients in the subcohort are identified by sparse linear tests, where the coordinate weights (coefficients) for the test hyperplane are given in Figure 3(c,d).
+Interestingly, this particular subcohort of patients shows significant methylation increase in BRCA2 and MSH6.
 Moreover, nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed when compared to the average non-metastatic patient in the dataset.
-Therefore, LXR-inverse agonists (see for example CVS+ (19)) may prove to be a useful therapeutic target for this subcohort of patients, if indicated by further clinical evidence.
+Therefore, LXR-inverse agonists (see for example lxr-ant) may prove to be a useful therapeutic target for this subcohort of patients, if indicated by further clinical evidence.
 (a) The nuclear receptors in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values.
@@
 (b) The methylation percentages in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values.
-In particular, for patients in this subcohort, BRCA2 and MSH6 were methylated more, both are related to DNA repair, and errors can lead to breast cancer progression SBS+ (19).
+In particular, for patients in this subcohort, BRCA2 and MSH6 were methylated more, both are related to DNA repair, and errors can lead to breast cancer progression brca2.
 Statistically insignificant bars are not shown.
@@
 In particular, in the above subcohort of patients, BRCA2 and MSH6 have significantly higher methylation percentages than the entire cohort of the patients in METABRIC dataset; and at the same time they have a significantly higher expression of NR1H2 receptors responsible for lipid metabolism.
-2.2.3 Identifying "good" subcohorts
-Recall that, searching over every pair of potential subsampled inputs and subcohorts is a computationally intractable problem, since there are an exponential number of such pairs.
-For the small dataset with 2000 patients, and subsamples of size 1000 patients, we may need to search through $2^{1000}$ runs of Algorithm 1!
-However, we can train a neural net to predict when we will have subcohorts of patients that are likely to lead to statistically significant associations, i.e., the pairs of subcohorts of patients will have enough common patients to lead to narrow confidence intervals.
-We can then prioritize our confidence interval calculations for such subcohorts.
-We implemented this prioritization using a simple single hidden layer neural net.
-We plot the quality of the neural net model, i.e., the errors in Figure 4 in the supplement.
-Even with a single hidden layer neural network, we see reasonable error margins.
-Thus prioritization is an easy problem to learn.
 5 Conclusion and limitations
-In conclusion, we have demonstrated a method to design tests that identify homogeneous subcohorts of patients.
-This was done by a non-trivial extension of the Goemans-Williamson algorithm GW (95), which may be independently interesting.
-Furthermore, such subcohorts may be used to design trials that help identify new treatments.
-In this paper, we have used the identified subcohorts to link epigenetic changes to changes in nuclear receptor expression levels.
-However, we do not have the resources to verify our predicted associations and any clinical predictions in an actual experimental setting and we hope this work inspires physicians to pursue that.
-4 Acknowledgments
-The author is grateful to Dr Siddhartha Mukherjee, and members of the Langelier lab (Dr.
-Natasha Spottiswoode, Dr Van Phan, and Dr Chaz Langlier) for mentioning the underlying clinical problem to him in two different patient-care settings – in patients with AML and sepsis, respectively.
-The author is also grateful to Dr Jacob Glass for pointing out the public datasets at the cbio portal, and to Professors Subhash Khot, Anupam Gupta and Cliff Stein for helpful comments.
+In conclusion, we have demonstrated a method to design tests that identify homogeneous subgroups (subcohorts) of patients.
+This was done by translating the problem to a MAX-CUT like formulation (as opposed to a rule-based decision tree like formulation), which admits a nearly optimal efficient approximation algorithm via an extension of the Goemans-Williamson algorithm gw.
+The latter may be independently interesting.
+Furthermore, the identified subcohorts and accompanying tests could help broaden the set of patients for existing treatments (for e.g.
+breast cancer patients and LXR inverse agnosts).
+However, we do not have the resources to verify our clinical predictions and we hope our algorithm (or similar idea) can be adapted by oncologists to pursue it further.
+6 Proof of Theorem 3.2
+Proof:
+We need to show that if the SDP solution has value OPT then the rounded solution has value at least $\alpha\mathrm{OPT}$, for some constant $\alpha$ (to be determined).
+Hence it suffices to upper-bound the ratio of each term in the in the SDP solution objective by the corresponding term in the objective for the expectation of the rounded solution.
+In other words, we need to upper-bound:
+$r:=\frac{2-2\langle v_{u},v_{u^{\prime}}\rangle-2\langle v_{u_{0}},v_{u^{\prime}}\rangle+2\langle v_{u},v_{u_{0}}\rangle}{\mathbb{E}\left[(z_{u}-z_{u^{\prime}})^{2}+(z_{u_{0}}-z_{u^{\prime}})^{2}-(z_{u}-z_{u_{0}})^{2}\right]},$ (13)
+where $z_{u},z_{u^{\prime}}$ and $z_{u_{0}}$ are obtained upon rounding $v_{u},v_{u^{\prime}}$ and $v_{u_{0}}$ using the Gaussian rounding procedure defined in Algorithm 1.
+Now imagine the three unit vectors $v_{u},v_{u^{\prime}}$ and $v_{u_{0}}$ as lying on the unit circle, and let $x\in[0,\pi]$ be the angle between $u_{0},u$ and $x^{\prime}\in[0,\pi]$ be the angle between $u_{0},u^{\prime}$.
+Next, consider each term in the denominator, $\mathbb{E}\left[(z_{u_{0}}-z_{u^{\prime}})^{2}\right]$ equals the probability that the random chosen Gaussian – that is equivalent to choosing a uniformly random point on the unit circle (with high probability) – is at an acute angle with $u_{0}$ and an obtuse angle with $u^{\prime}$ or vice versa.
+Thus,
+$\mathbb{E}\left[(z_{u_{0}}-z_{u^{\prime}})^{2}\right]=4\cdot\frac{x^{\prime}}{\pi}.$ (14)
+Similarly, one has:
+$\mathbb{E}\left[(z_{u_{0}}-z_{u})^{2}\right]$ $=$ $4\cdot\frac{x}{\pi}$ (15)
+Two cases arise:
+$x+x^{\prime}\leq\pi:\ \mathbb{E}\left[(z_{u^{\prime}}-z_{u})^{2}\right]$ $=$ $4\cdot\frac{x+x^{\prime}}{\pi}$ (16)
+$x+x^{\prime}\geq\pi:\ \mathbb{E}\left[(z_{u^{\prime}}-z_{u})^{2}\right]$ $=$ $4\cdot\frac{2\pi-x-x^{\prime}}{\pi}.$ (17)
+Furthermore, the numerator equals:
+$2-2\langle v_{u},v_{u^{\prime}}\rangle-2\langle v_{u_{0}},v_{u^{\prime}}\rangle+2\langle v_{u},v_{u_{0}}\rangle=2-2\cos(\theta)-2\cos(x^{\prime})+2\cos(x),$ (18)
+where we have used $\cos(\pi-x)=\cos(x)$ and $\theta$ is either $x+x^{\prime}$ (for $x+x^{\prime}\leq\pi$), or $\theta$ is either $2\pi-x-x^{\prime}$ (for $x+x^{\prime}\geq\pi$).
+Thus the ratio $r$ in Equation 13 equals: [^5] Just to compare, for the well-known rounding in [gw], one finds the corresponding ratio to be: $r_{GW}=\frac{1-\cos(x)}{x}$, and the maximum is well-known to be at $x\simeq 134\mathrm{deg}$.
+$x+x^{\prime}\leq\pi:\ r(x,x^{\prime})=\frac{\pi}{4}\cdot\frac{2-2\cos(x+x^{\prime})-2\cos(x^{\prime})+2\cos(x)}{2x^{\prime}}$ (19)
+$x+x^{\prime}\geq\pi:\ r(x,x^{\prime})=\frac{\pi}{4}\cdot\frac{2-2\cos(2\pi-x-x^{\prime})-2\cos(x^{\prime})+2\cos(x)}{2\pi-2x}.$ (20)
+Note that whether $x^{\prime}\to 0$ in Equation 19 or $x\to\pi$ in Equation 20, the denominator goes to 0 but the numerator goes to 0 as well.
+Moreover, since $\cos(x)$ is even, the numerator goes to 0 as a quadratic in both cases, so the limit is 0.
+Thus 0 or $\pi$ are likely not the maximum points.
+For $x+x^{\prime}\leq\pi$: Taking the partial derivative with respect to $x$, one gets: $\sin(x+x^{\prime})-\sin(x)=0$.
+Thus either (1) $x+x^{\prime}=x$, so $x^{\prime}=0$; or (2) $x+x^{\prime}=\pi-x$, so $2x=\pi-x^{\prime}$.
+However, we have ruled out (1).
+Substituting (2) into Equation 19, we get (at the extremum):
+$r(x)$ $=$ $\frac{\pi}{4}\cdot\frac{1-\cos(x+x^{\prime})-\cos(x^{\prime})+\cos(x)}{x^{\prime}}$ (21)
+$=$ $\frac{\pi}{4}\cdot\frac{1-\cos(\pi/2+x^{\prime}/2)+\cos(\pi/2-x^{\prime}/2)-\cos(x^{\prime})}{2x^{\prime}}.$ (22)
+Plotting the single variable function above, shows that the maximum value of $r$ is $1.22$ and is reached at x=1.86x=1.86rads.
+For $x+x^{\prime}\geq\pi$: Taking the partial derivative with respect to $x^{\prime}$ and equating with 0 shows: $\sin(x^{\prime})-\sin(x+x^{\prime}-\pi)=0$ at the extremum.
+Thus either (1) $x^{\prime}=x+x^{\prime}-\pi$, so $x=\pi$; or (2) $\pi-x^{\prime}=x+x^{\prime}-\pi$, so $x^{\prime}=\pi-x/2$.
+However, we have ruled out (1).
+Substituting with (2) into Equation 20, we get (at the extremum):
+$r(x)=\frac{\pi}{4}\cdot\frac{1-\cos(\pi-x/2)-\cos(\pi-x/2)+\cos(x)}{\pi-x}=\frac{\pi}{4}\cdot\frac{1-2\cos(\pi-x/2)+\cos(x)}{\pi-x}.$ (23)
+Plotting the single variable function above, shows that the maximum value of $r$ is $1.22$ and is reached at x=1.28x=1.28rads.
+[…]
+7 Proof of Theorem 3.3
+Proof:
+Without loss of generality assume that the variable $z_{u_{0}}$ in Equation 1 is +1+1 (if not one can flip all the signs).
+Next, note that each term of the objective of the optimization problem in Equation 1 is either +1+1 or 1-1 (for the $\pm 1$ solution).
+The total value of the cut is assumed to be $\kappa\cdot|U||V|$.
+Therefore, if $x$ and $y$ denote the fraction of $U$ and $V$ vertices on the subcohort side of the cut, then we have by comparing the leading term in the objective with the cut value above:
+$2x(1-y)\cdot|U||V|\geq\kappa\cdot|U||V|$ (24)
+However, the cut we compute using SDP rounding is not optimal but only close to it, i.e.,
+$2x(1-y)\cdot|U||V|\geq\alpha\kappa\cdot|U||V|,$ (25)
+with $\alpha\simeq 0.82$ (cf. Theorem 3.2).
+Finally, the specificity is the ratio of $U$ points on selected side of the cut with all $U$ and $V$ points on the subcohort side of the cut, i.e., specificity is given by the ratio of $U$ vertices over all vertices on the subcohort side of the cut, i.e.,
+$\frac{x|U|}{y|V|+x|U|}$ (26)
+In order to lower bound the ratio, we simply need to compute its minimum over $x,y\in(0,1)$ under the constraint in Equation 25. Substituting $y$ with 1y1-y proves the result. $\square$
+One may think of $\kappa$ as the maximum fraction of edges that can be in the cut given the sparsity constraints.
+Thus $\kappa$ is the upper bound on sensitivity, while the value of the optimum in Equation 9 is a lower bound on the sensitivity.
+Thus we plot the objective in Equation 9 in Figure 4 for various values of $\kappa$ and $\frac{|V|}{|U|}$, to illustrate the lower bound on specificity.
+(a) Heatmap depicting lower bound on specificity as a function of $\kappa$ (x-axis) and $\frac{|V|}{|U|}$ (y-axis)
+(b) Lowerbound on specificity vs $\kappa$ (max. possible sensitivity)
+Figure 4: Lower bound on specificity as a function of $\kappa$ and $\frac{|V|}{|U|}$, based on Theorem $3.$
+8 Proof of Theorem 3.4
+Proof:
+A straight forward application of Lemma 8.1 (also known as Grothendieck’s identity) implies that the RHS of the constraints in Equations 6 and 8 equals $\frac{2\theta}{\pi}$ and $\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ respectively, in expectation.
+Thus the constraints of our SDP relaxation are not violated excessively.
+Hence the proof follows.
+$\square$
+Lemma 8.1
+Given a $2$-dimensional Gaussian $(X,Y)$ with mean zero and covariance matrix $\begin{bmatrix}\delta_{x}&\rho\\ \rho&\delta_{y}\end{bmatrix}$, we have:
+$\mathbb{E}[\mathrm{sign}(X)\mathrm{sign}(Y)]=1-2\mathbb{P}\left(\frac{Z_{1}}{Z_{2}}\leq t\right)=\frac{2}{\pi}\arcsin\left(\frac{\rho}{\sqrt{\delta_{x}\delta_{y}}}\right),$ (27)
+where $Z_{1}$ and $Z_{2}$ are independent mean zero Gaussians with covariance $\rho$ and variance $1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}$ and $\delta_{y}$ respectively.
+Proof:
+Note that, $Z:=\frac{X}{\sqrt{\delta_{x}}}-\frac{\rho Y}{\sqrt{\delta_{x}}\delta_{y}}$ is a mean zero Gaussian independent of $Y$, and has variance $1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}$.
+$\mathbb{P}\left(XY\geq 0\right)$ $=$ $\mathbb{P}\left(ZY\geq\frac{\rho}{\sqrt{\delta_{x}}\delta_{y}}Y^{2}\right)$ (28)
+$=$ $\mathbb{P}\left(\frac{Z}{\sqrt{1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}}}\geq\frac{\rho Y}{\sqrt{\delta_{x}}\delta_{y}\sqrt{1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}}}\right)$ (29)
+$=$ $\mathbb{P}\left(\frac{\tilde{Z}}{\tilde{Y}}\geq\frac{\rho}{\sqrt{\delta_{x}\delta_{y}}\sqrt{1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}}}\right),$ (30)
+where $\tilde{Y}$ and $\tilde{Z}$ are independent standard Gaussians, and thus their ratio is the Cauchy distribution. Therefore, the last probability above can be calculated as:
+$\mathbb{P}\left(XY\geq 0\right)$ $=$ $\frac{1}{\pi}\arctan\left(\frac{\rho}{\sqrt{\delta_{x}\delta_{y}}\sqrt{1+\frac{\rho^{2}}{\delta_{x}\delta_{y}}}}\right)$ (31)
+$=$ $\frac{1}{\pi}\arcsin\left(\frac{\rho}{\sqrt{\delta_{x}\delta_{y}}}\right).$ (32)
+This completes the proof of Lemma 8.1. $\square$
+9 Comparison with PRIM (continued from Subsection 4.1)
+In this section, we compare our algorithm’s performance against local search based algorithms (PRIM based subgroup discovery).
+As a baseline, we implemented a version of the PRIM subgroup discovery algorithm [ff].
+PRIM is a local search algorithm, where the subgroups are defined by axis-parallel hyper-rectangles.
+The algorithm work as follows: One starts with a hyper-rectangle covering the entire dataset, and peels off a small portion of the rectangle from one or more dimensions (sides) in each iteration of the algorithm.
+In each iteration, one computes and stores the specificity for the subgroup defined by the points within the rectangle and after trying a sufficient number of rectangles one returns the optimal subgroup (with highest specificity) seen thus far.
+In this way the algorithm gradually trades-off sensitivity for increased specificity.
+In our case, we additionally want to ensure a sparsity constraint, and then compute the specificity for a fixed sensitivity in order for comparison with our algorithm.
+So we make the following modifications:
+  • We enforce our sparsity constraint by using rectangles of dimension at most $k$ that are chosen as random subsets of size $k$ from $\{1,...,50\}$.
+The reason for using $50$ is that in our case the subsets correspond to subsets of the 50 possible nuclear receptors.
+  • We require that the generated rectangle have at least $\sigma$ fraction of points that correspond to patients with metastatic disease ($\sigma$ was chosen to be either 0.1 or 0.3 for comparison).
+Thus ensuring a minimal sensitivity for the generated subgroups.
+Finally, we capped our search by generating 400K axis-parallel rectangles. It required about $5\times$ the computational time needed compared to Algorithm 1.
+Baseline (PRIM) vs Experiment (MAX-CUT) specificity: The results for the rule-based local search PRIM algorithm are shown in Figure 5.
+Figure 5: Plot specificity vs k (sparsity) as sensitivity is held above a lower bound for the variant of PRIM algorithm used for comparison.
+The dashed Red line (in both figures) denotes the specificity for the MAX-CUT algorithm, from Figure 1, for the given sensitivities ($0.3$ and $0.1$), when the number of genes used is 15 (L figure) and 45 (R figure).
+In both cases, the PRIM local search heuristic fails to approach the specificity of Algorithm 1.
+Therefore, the MAX-CUT based linear separator approach shows better specificity than the rule-based PRIM algorithm for a given sensitivity.
+This is not unexpected as Algorithm 1 is not a simple local search confined to an axis-parallel hyper-rectangle, but a more sophisticated SDP rounding based algorithm that computes a general (non-axis-aligned) sparse linear separator.